# ERABAL: Enhancing Role-Playing Agents through Boundary-Aware Learning

### v1

@@ Abstract
Role-playing is an emerging application in the field of Human-Computer Interaction (HCI), primarily implemented through the alignment training of a large language model (LLM) with assigned characters. Despite significant progress, role-playing agents (RPLAs) still struggle with maintaining role-consistency across conversations, particularly when confronted with boundary queries subtly related to character attributes. In this paper, we present Erabal, a framework aimed at enhancing RPLAs’ role-playing capabilities through boundary-aware learning. Erabal encompasses a generation pipeline for role-specific dialogues and a concomitant methodology for alignment training. Through comprehensive evaluations, we demonstrate that Erabal is both efficient and effective. By training with significantly fewer dialogues than those used in leading approaches, Erabal achieves notable improvements across WikiRoleEval, CharacterEval, and the role-playing subset of MT-Bench compared to the generalist baseline models. Our code and datasets will be made publicly available to support further research.
@@ 1 Introduction
Role-playing is an emerging research field Chen et al. (2024); Tseng et al. (2024) in which AI agents engage in dialogues with humans by simulating the linguistic style or behavioral traits of the assigned characters (Ouyang et al., 2022; Yi et al., 2024). As large language models (LLMs) have shown great potential for achieving human-level intelligence, their use in role-playing agents (RPLAs) (Shanahan et al., 2023; Mao et al., 2023; Zhou et al., 2023; Li et al., 2023a; Park et al., 2023) has garnered increasing attention, leading to remarkable progress recently. Nonetheless, RPLAs still struggle to maintain role consistency across multi-turn dialogues, especially when confronted with queries that are subtly related to character attributes but not explicitly mentioned within the prompts, resulting in an out-of-character (OOC) issue.
@@
Figure 1: Illustration of the boundary-aware dialogue. Role-specific boundary-aware queries and responses are near the edge, while ordinary queries and responses are at the center. Unknown queries outside the character’s scope are met with rejection responses. Both positive and negative responses are needed in this work.
@@
There have been various efforts towards addressing this challenge (Shao et al., 2023; Yu et al., 2024; Lu et al., 2024). A notable approach among them, proposed by Ditto (Lu et al., 2024), constructs dialogues with “inter-role contrastive queries". These queries are answerable for certain roles but unknown to others as they go beyond their era or occupation. Through alignment training with such queries and their rejection responses, Ditto proves its effectiveness in tackling the out-of-character issue (Shao et al., 2023). However, as contrastive queries derived from different roles are indeed coarse-grained, there is still room for improvement by constructing training data through a more fine-grained strategy.
In this paper, we propose Erabal, a boundary-aware learning framework targeted at enhancing RPLAs’ role-playing capabilities. Erabal encompasses a boundary-aware dialogue generation pipeline and a concomitant methodology for both supervised fine-tuning and post-training. As Figure 1 illustrates, a boundary-aware dialogue here is expected to meet the following conditions: (1) It contains queries that are highly role-specific and cannot be simply answered by rephrasing the role description within the prompt, nor can they be easily dismissed with rejection answers. (2) Among all suitable answers to a boundary query, the expected one should comprehensively reflect the character attributes where possible. The boundary-aware dialogues delineate the nuanced boundaries of a role, compelling an LLM to fully utilize its inherent character knowledge to better serve as an RPLA.
@@
Erabal is designed to produce such dialogues. It introduces an automatic data generation pipeline with four modules: a dialogue planner, a topic manager, a boundary query generator, and a response generator. During data construction, the boundary query generator first selects a snippet from an external character attribute resource (e.g., Wikipedia111www.wikipedia.org) and then tampers it as a counterfactual prerequisite to prompt an LLM (e.g., GPT-4) to generate a role-specific query. The original snippet is further passed to the response generator, and is leveraged to identify the subtle clues and generate a factual response that is consistent with the assigned character. Boundary-aware dialogues consisting of such question-answer turns are used as training data in the supervised fine-tuning stage. To depict fine-grained boundaries for roles, we also introduce the boundary-aware preference optimization (BPO) method. Following DPO (Rafailov et al., 2023), BPO also leverages a contrastive learning instead of RL learning paradigm based on the synthesized sample pairs. We use the factual response as the positive sample and generate a counterfactual response as the negative sample by providing the response generator with the tampered snippet. The factual and counterfactual responses, along with the dialogue context, are paired to form a training sample, aswill be used during the BPO post-training stage.
Models Boundary Role-Consistency MT-Bench Role-play
@@
Table 1: Experimental results of Baichuan2-13B-Chat and Mistral-7B-Instruct-v0.3 in boundary scenario and MT-Bench role-playing evaluation. Erabal achieves an average of 0.17 (out of 1.0) and 0.38 (out of 10) absolute improvements over the generalist baselines.
To serve as a proof-of-concept, we collect 90 popular roles and 500 boundary-aware dialogues as the test set through Erabal. We then assess existing generalist models and their variants through the role consistency evaluation metric proposed in Ditto. As shown in Table 1, generalist models such as Baichuan2-Chat Yang et al. (2023) and Mistral-Instruct Jiang et al. (2023a) perform sub-optimally, although those enhanced with contrastive queries exhibit superior performances. As a comparison, variants enhanced with Erabal demonstrate the best performance in this scenario. More details will be detailed in Section 3.1. We also conduct comprehensive experiments across WikiRoleEval (Lu et al., 2024), CharacterEval (Tu et al., 2024), and the role-playing subset of MT-Bench (Zheng et al., 2023). In this case, we additionally collect a training dataset containing 100 roles and 8,369 training dialogues, where the roles are identical to those in RoleBench [^2] RoleBench is one of the largest open-source role-playing datasets. (Wang et al., 2023c) for a fair comparison. Additionally, we collect a variant of our training dataset named Erabal ditto where the boundary-aware question-answer turns are replaced by contrastive queries and rejection responses sampled following (Lu et al., 2024). Experimental results demonstrate that Erabal is both efficient and effective: using only 10% of the training dialogues than those used in RoleBench, Erabal achieves significant improvements over the generalist baselines across the three benchmarks, demonstrating a significant advantage of our method over previous approaches.
Our contributions are as follows: (1) We introduce the boundary-aware learning method and demonstrate its effectiveness in building role-playing agents. (2) We propose Erabal, a framework encompassing a multi-module data production pipeline and the concomitant methodology for supervised fine-tuning and post-training. (3) Extensive experiments show that our method significantly improves performance across existing role-playing benchmarks, particularly enhancing metrics closely related to the role-playing ability.
@@ 2 Boundary-aware Dialogue Construction
Figure 2 shows the overall data generation process. To construct a boundary-aware dialogue, we begin with the dialogue planner and topic manager. The two modules generate instructions at the start of each turn, determining whether to generate an “ordinary” or “boundary” query and specifying the conversational topic. Here we introduce ordinary questions to ensure the constructed dialogue remains coherent, thereby making RPLAs with such data practical for real applications. The instructions are then passed to the query generator and the response generator, which are responsible for generating specific question-answer pairs based on external character attributes. In the following section, we will first describe each component, then introduce the quality control strategies, and finally provide an overall statistic of the constructed dataset.
@@ 2.1 Main Components
The response generator is responsible for generating responses based on the given character attributes, conversational context, and snippets. Note that snippets are passed to this module as references only when the boundary instruction are specified. The response generator is expected to produce both factual and counterfactual responses corresponding to the original and tampered snippets respectively. In practice, even if we provide an original snippet, the response generator may still produce a counterfactual response, and vice versa. Therefore, we additionally introduce an information verifier to perform secondary verification to ensure the responses meet our demands. Detailed prompts are shown in Table 13 and Table 14.
@@ 2.2 Quality Control
To ensure the dialogues generated by Erabal meet our goal, we recruited 3 annotators with graduate degrees for \$20 an hour to assess their quality based on the questions listed in Table 2. The four questions range from simple to complex, and an ideal dialogue should meet all of them simultaneously. According to the statistics, about 26% of the dialogues fail to address the most complex questions. Indeed, this is already not bad, as it is challenging to compel LLMs to formulate boundary queries and embed counterfactual traps within them. Besides, the failed dialogues do not hinder the training of RPLAs, as they simply revert to ordinary role-specific dialogues.
@@ 3.1 Boundary Scenario Evaluation
To examine if boundary-aware training can mitigate OOC issues, we adopt the role consistency metric proposed in Ditto for evaluation. Note that knowledge accuracy evaluation is unsuitable in this scenario, as the queries and responses are too ordinary to reflect characters’ boundaries. Moreover, the straightforward presentation of role-specific knowledge makes it difficult to align the evaluation with the characters’ self-awareness. We assess a response as a multiple-choice problem: it is compared against four character options—one ground truth and three randomly sampled from the character set. GPT-4, serving as a judge, is tasked with identifying the option that best matches the response. The prompt for evaluation can be found in Table 17.
@@ 3.2 Comprehensive Evaluations
We further examine if boundary-aware learning can enhance the general role-playing capabilities. Here we adopt WikiRoleEval, CharacterEval, and the role-playing subset of MT-Bench for evaluation.
@@ 3.2.1 WikiRoleEval
WikiRoleEval is an evaluation benchmark that focuses on three dimensions of role-playing capabilities: role consistency, knowledge accuracy, and unknown question rejection. The first two dimensions respectively assess the RPLA’s ability on human likeness and knowledge, whereas the last dimension focuses on evaluating the RPLA’s accuracy in rejecting out-of-character questions. Detailed descriptions can be found in Appendix B.2.
@@ 3.2.2 CharacterEval
CharacterEval includes a set of 12 metrics designed to comprehensively assess RPLAs, which are divided into three categories: role consistency, conversational ability, and role-playing attractiveness. It emphasizes more on evaluating the immersive conversational experience, rather than addressing other issues. Detailed descriptions can also be found in Appendix B.3.
@@ 3.2.3 MT-Bench
MT-Bench has well-designed questions spanning eight categories, including a subset of role-playing. We use GPT-4 to evaluate the RPLA’s responses through 2-turn evaluation dialogues.
@@ 4.1 Experimental Setups
For the boundary scenario evaluation, we use Baichuan2, Qwen Bai et al. (2023), Mistral, and LLaMA2 (Touvron et al., 2023) as base models. We assess four variants of these models using the following configurations: (1) without fine-tuning (i.e., the generalist chat model), (2) fine-tuning with RoleBench, (3) fine-tuning with contrastive queries sampled following Ditto’s strategy, and (4) fine-tuning with Erabal’s 100 role training dataset.
For comprehensive evaluations, we select Baichuan2 and Mistral as base models and conduct supervised fine-tuning followed by BPO post-training based on Erabal’s entire dataset.
@@ 4.3 Baselines
(1) Generalist Baselines. OpenChat-3.5 (Wang et al., 2023a), GPT-3.5-Turbo (OpenAI, 2022), and GPT-4 (OpenAI, 2023) represent state-of-the-art performance in general-purpose conversational models. Claude-2.1 (Anthropic, 2023) and Wenxin-4333https://yiyan.baidu.com/ emerge as formidable competitors, excelling in the English and Chinese languages, respectively.
(2) Dedicated Baselines. CharacterGLM (Zhou et al., 2023), Character-LLM (Park et al., 2023) are two open-source dedicated models optimized for role-playing. Xingchen is a close-source role-playing platform capable of creating any character with a given persona. Ditto is a closed-source model trained through the self-alignment method, based on the Qwen w/o role-play model.
@@ 4.4.1 Boundary Scenario Evaluation Results
Figure 3 shows the results of the baselines and variants of Erabal. The four generalist chat models, without exception, perform sub-optimally in boundary scenarios and are still far from achieving satisfactory results in terms of role consistency performance. This indicates that LLMs without role-playing enhancement struggle to maintain their persona across multi-turn dialogues, experiencing an out-of-character (OOC) issue. Besides, we do not observe a clear advantage of RPLAs trained with the general role-playing dataset (i.e., RoleBench) over the baselines. It suggests that acquiring concise boundary awareness may significantly differ from learning general role-playing abilities. While the characters’ linguistic styles or behavior patterns can be easily mimicked, it remains challenging for RPLAs to fully grasp the complete picture of the characters.
We also conduct experiments with the inter-role contrastive query sampling strategy based on Erabal’s 100 roles training set. RPLAs trained with such data demonstrate significant improvements. However, due to the coarse-grained nature of the sampling strategy, they still perform unsatisfactorily. Ultimately, RPLAs trained with Erabal effectively improve the accuracy score by an average absolute of 20.4 points, culminating in superior performance across all baselines.
@@ 4.4.2 Comprehensive Evaluation Results
Table 4 shows the evaluation results on WikiRoleEval. We observe that different categories of models exhibit significant variations in performance across the three dimensions. Generalist baseline models, including OpenChat-3.5, Claude-2.1, Wenxin-4, GPT-3.5-Turbo, and GPT-4 series, perform better on knowledge accuracy and unknown question rejection tasks compared to role consistency. Knowledge accuracy is a task highly related to the model size, while unknown question rejection is relatively coarse-grained, which may be easy for LLMs to handle. Role consistency, which implicitly involves fine-grained boundary-aware assessments of an RPLA, presents a greater challenge to generalist models.
@@
Experimental results on the role-playing subset of MT-Bench in Table 4 show a consistent trend as WikiRoleEval. However, the performance enhancements are limited, especially between the supervised fine-tuning and BPO post-training stages. We conjecture that this is because the role-play task in MT-Bench focuses more on functional roles rather than human-like characters, making it less sensitive to our boundary-aware learning approach.
@@
Table 5 exhibits the evaluation results on CharacterEval. It can be observed that Erabal significantly improves the baseline models’ performances on most metrics, especially on persona behavior (PB), role consistency (Cons.), and human likeness (HL). These metrics have high correlation to RPLAs’ general role-playing capabilities. BPO further improves the aforementioned three metrics, demonstrating its effectiveness.

### v2

@@ Abstract
Role-playing is an emerging application in the field of Human-Computer Interaction (HCI), primarily achieved through the alignment training of large language models (LLMs) with various characters. Despite significant progress, role-playing agents (RPLAs) still encounter challenges in maintaining character trait and factual consistency during conversations, particularly when confronted with boundary queries subtly related to character attributes. In this paper, we present Erabal, a boundary-aware learning framework designed to enhance the role-playing capabilities of RPLAs. Erabal encompasses a data generation pipeline and a concomitant alignment training methodology, aimed at generating and learning from boundary samples respectively. Through comprehensive evaluations, Erabal demonstrates significant improvements across WikiRoleEval, CharacterEval, and the role-playing subset of MT-Bench, outperforming generalist baseline models while requiring considerably fewer dialogues. Our code and datasets will be made publicly available to support further research.
@@ 1 Introduction
Role-playing is an emerging research field Chen et al. (2024); Tseng et al. (2024) where AI agents engage in dialogues with humans by simulating the linguistic styles or behavioral traits of the assigned characters (Ouyang et al., 2022; Yi et al., 2024). As large language models (LLMs) have shown great potential for achieving human-like intelligence, their use in role-playing agents (RPLAs) (Shanahan et al., 2023; Mao et al., 2023; Zhou et al., 2023; Li et al., 2023a; Park et al., 2023) has garnered increasing attention, leading to remarkable progress recently. Nonetheless, RPLAs still struggle to maintain consistency during human-computer dialogues, facing issues such as character trait contradictions and factual errors. Figure 1 shows such a case, where the RPLA takes on the character of Harry Potter and confronts with the query about “magical creatures". Models that have not been enhanced by boundary-aware learning are prone to falling into traps, resulting in out-of-character (OOC) issues.
@@
Figure 1: Illustration of boundary-aware learning. The circular areas represent the cognitive scope of the characters. High-value queries (marked in blue) are located within and near the boundaries, while ordinary queries (marked in yellow) are closer to the center. Unknown queries (marked in red), which should be dismissed with rejection responses, fall outside the circles. Both positive and negative responses are generated in our work.
@@
There have been various efforts towards addressing this challenge (Shao et al., 2023; Yu et al., 2024; Lu et al., 2024). A notable approach among them, proposed by Ditto(Lu et al., 2024), constructs training dialogues with inter-role contrastive queries. These queries are devised to be answerable for specific roles but fall outside the cognitive scopes of others, as they transcend their era or occupation. By drawing rough boundaries for characters through alignment training, Ditto demonstrates its effectiveness in improving RPLAs’ role-playing capabilities. However, since this method is relatively coarse-grained, there is still room for improvement by training through a more fine-grained strategy.
@@
In this paper, we propose Erabal, a boundary-aware learning framework aimed at enhancing RPLAs’ role-playing capabilities. Erabal introduces an automatic data generation pipeline consisting of four modules: a dialogue planner, a topic manager, a query generator, and a response generator. During training data construction, the query generator first selects a snippet from a comprehensive external character attribute resource (e.g., Wikipedia111www.wikipedia.org) and then tampers it as a counterfactual prerequisite to prompt an LLM (e.g., GPT-4) to generate a role-specific query. The character attribute resource is not visible to LLMs during training, and a short role description is provided instead. Such queries are boundary-aware since they are highly role-specific and cannot be easily answered through the general knowledge in LLMs. The original snippet is further passed to the response generator, and is leveraged to generate a factual response that is consistent with the assigned character. Dialogues consisting of such boundary-aware query-response pairs are used as supervised fine-tuning data. To depict fine-grained boundaries for roles, we also introduce a post-training stage. We pair each aforementioned factual response with a counterfactual response as the negative sample by providing the response generator with the tampered snippet. The pairs, along with their dialogue context, are utilized by the direct preference optimization (DPO) (Rafailov et al., 2023) method.
Models Boundary Role Consistency MT-Bench Role-play
@@
Table 1: Experimental results of Baichuan2-13B-Chat and Mistral-7B-Instruct-v0.3 in boundary scenario evaluation (Section 3.1) and MT-Bench role-playing evaluation. Erabal achieves an average of 0.17 and 0.38 absolute improvements over the generalist baselines.
To serve as a proof of concept, we first collect 90 popular roles and 500 boundary-aware dialogues as the test set using Erabal. We evaluate generalist models and their variants through the role consistency evaluation proposed in Ditto. As shown in Table 1, generalist models such as Baichuan2-Chat Yang et al. (2023) and Mistral-Instruct Jiang et al. (2023a) perform sub-optimally, while those enhanced with contrastive queries indeed demonstrate superior performance. In comparison, variants enhanced with Erabal demonstrate the best performance (will be detailed in Section 3.1). We also conduct comprehensive experiments across WikiRoleEval (Lu et al., 2024), CharacterEval (Tu et al., 2024), and the role-playing subset of MT-Bench (Zheng et al., 2023). In this case, we additionally collect a training dataset containing 190 roles and 16,932 dialogues, which includes 100 roles that are identical to those in RoleBench [^2] RoleBench has only 5 Chinese roles; we collect an additional 90 Chinese roles for the comprehensive evaluation. (Wang et al., 2023c). Experimental results demonstrate that Erabal is both efficient and effective: by using only 10% of the training dialogues compared to those in RoleBench ( 168,093), Erabal achieves significant improvements over the generalist baselines across the three benchmarks, highlighting a clear advantage of our method over previous approaches.
Our contributions are as follows: (1) We introduce the boundary-aware learning method and demonstrate its effectiveness in developing role-playing agents. (2) We propose Erabal, a framework encompassing a multi-module data production pipeline and the concomitant methodology for supervised fine-tuning and post-training. (3) Extensive experiments show that our method significantly enhances performance across existing role-playing benchmarks, particularly enhancing metrics closely related to role-playing abilities.
@@ 2 Boundary-aware Dialogue Construction
Figure 2 shows the overall data generation process. To construct a boundary-aware dialogue, we begin with the dialogue planner and topic manager. These two modules generate instructions at the start of each turn, determining the question type and specifying the conversational topic. The question type can be either “ordinary” or “boundary”. Ordinary questions are introduced to ensure that the constructed dialogue remains coherent, which is essential for maintaining contextual relevance during conversations. The instructions are then passed to the query generator and the response generator, which are responsible for generating specific question-answer pairs based on external character attributes. The external character attributes are replaced with a brief description in the final training data. In the following section, we will first describe each component, then introduce the quality control strategies, and finally provide an overall statistic of the constructed dataset.
@@ 2.1 Main Components
The response generator is responsible for generating responses based on the given character attributes, conversational context, and snippets. Note that snippets are passed to this module as references only when the boundary instruction are specified. The response generator is expected to produce both factual and counterfactual responses corresponding to the original and tampered snippets respectively. In practice, even if we provide an original snippet, the response generator may still produce a counterfactual response, and vice versa. Therefore, we additionally introduce an information verifier to perform secondary verification to ensure the responses meet our demands. Detailed prompts are listed in Table 16 and Table 17.
@@ 2.2 Quality Control
To ensure the dialogues generated by Erabal meet our goal, we recruited 3 annotators with graduate degrees for \$20 an hour to assess their quality based on the questions listed in Table 2. The four questions range from simple to complex, and an ideal dialogue should meet all of them simultaneously. According to the statistics, about 26% of the dialogues fail to address the most complex questions. We believe this is a reasonable outcome, given the challenge of getting LLMs to formulate boundary queries and embed counterfactual traps within them. Fortunately, the failed dialogues do not hinder the training of RPLAs, as they simply revert to ordinary role-specific dialogues. In addition, we conducted ablation experiments on various modules. The results show that key components of Erabal, such as the topic manager and dialogue planner, are essential for enhancing data diversity. Please refer to Appendix B.4 for more details.
@@ 3.1 Boundary Scenario Evaluation
In this experiment, we adopt the role consistency metric proposed in Ditto for evaluation. It treats the evaluation as a multiple-choice problem: the model is expected to select a proper character from four for each response. GPT-4, serving as a judge, is tasked with identifying the option that best matches the response. We split the training and test sets according to roles to ensure that there is no character overlap between the training and testing datasets. Prompts for evaluation can be found in Table 20.
@@ 3.2 Comprehensive Evaluations
To investigate whether boundary-aware learning can enhance the general role-playing capabilities, we adopt WikiRoleEval, CharacterEval, and the role-playing subset of MT-Bench for evaluation in this experiment.
@@ 3.2.1 WikiRoleEval
WikiRoleEval (Lu et al., 2024) is an evaluation benchmark that focuses on three dimensions of role-playing capabilities: role consistency, knowledge accuracy, and unknown question rejection. The first two dimensions respectively assess the RPLA’s ability on human-likeness and knowledge, whereas the last dimension focuses on evaluating the RPLA’s accuracy in rejecting out-of-character questions. Detailed descriptions can be found in Appendix B.1.
@@ 3.2.2 CharacterEval
CharacterEval (Tu et al., 2024) includes a set of 12 metrics designed to comprehensively assess RPLAs, which are divided into three categories: role consistency, conversational ability, and role-playing attractiveness. It emphasizes more on evaluating the immersive conversational experience, rather than addressing other issues. Detailed descriptions can also be found in Appendix B.2.
@@ 3.2.3 MT-Bench
MT-Bench (Zheng et al., 2023) has well-designed questions spanning eight categories, including a subset of role-playing. We use GPT-4 to evaluate the RPLA’s responses through 2-turn evaluation dialogues.
@@ 4.1 Experimental Setups
For the boundary scenario evaluation, we use Baichuan2 Yang et al. (2023), Qwen Bai et al. (2023), Mistral Jiang et al. (2023a), and LLaMA2 (Touvron et al., 2023) as base models. We assess four variants of these models using the following configurations: (1) without fine-tuning (i.e., the generalist chat model), (2) fine-tuning with the RoleBench dataset, (3) fine-tuning with dialogues constructed using Ditto, and (4) fine-tuning with the Erabal 100 role training dataset.
For comprehensive evaluations, we select Baichuan2 and Mistral as base models and conduct supervised fine-tuning followed by DPO based on the entire Erabal all dataset.
@@ 4.3 Baselines
(1) Generalist Baselines. OpenChat-3.5 (Wang et al., 2023a), GPT-3.5-Turbo (OpenAI, 2022), and GPT-4 (OpenAI, 2023) represent state-of-the-art general-purpose conversational models. Claude-2.1 (Anthropic, 2023) and Wenxin-4333https://yiyan.baidu.com/ emerge as formidable competitors, excelling in the English and Chinese languages, respectively.
(2) Dedicated Baselines. CharacterGLM (Zhou et al., 2023), Character-LLM (Park et al., 2023) are two open-source dedicated models tuned for role-playing. Xingchen is a close-source role-playing platform capable of creating any character with a given persona. Ditto is a closed-source model trained through the self-alignment method, based on the Qwen w/o role-play model.
@@ 4.4.1 Boundary Scenario Evaluation Results
Figure 3 shows the results of the baselines and variants of Erabal. The four generalist chat models, without exception, perform sub-optimally in boundary scenarios and are still far from achieving satisfactory results in terms of role consistency performance. This indicates that LLMs without role-playing enhancement struggle to maintain persona consistency throughout multi-turn dialogues. Besides, we do not observe a clear advantage of RPLAs trained with the general role-playing dataset (i.e., RoleBench) over the baselines. It suggests that acquiring concise boundary awareness may significantly differ from learning general role-playing abilities. While the characters’ linguistic styles or behavior patterns can be easily mimicked, it remains challenging for RPLAs to fully grasp the complete picture of the characters.
We also conduct experiments with the inter-role contrastive query sampling strategy based on Erabal’s 100 roles training set. RPLAs trained with such data demonstrate significant improvements. However, due to the coarse-grained nature of the sampling strategy, there is still room for improvement. Ultimately, RPLAs trained with Erabal effectively improve the accuracy score by an average absolute of 20.4 points, culminating in superior performance across all baselines.
@@ 4.4.2 Comprehensive Evaluation Results
Table 4 shows the evaluation results on WikiRoleEval. We observe that different categories of models exhibit significant variations in performance across the three dimensions. Generalist baseline models, including OpenChat-3.5, Claude-2.1, Wenxin-4, GPT-3.5-Turbo, and GPT-4 series, perform better on knowledge accuracy and unknown question rejection tasks compared to role consistency. Knowledge accuracy is a task highly related to model size, while unknown question rejection is relatively coarse-grained, which may be easy for LLMs to handle. Role consistency, which implicitly involves fine-grained boundary-aware assessments of an RPLA, presents a greater challenge to generalist models.
@@
Experimental results on the role-playing subset of MT-Bench in Table 4 show a consistent trend as WikiRoleEval. However, the performance enhancements are limited, especially between the supervised fine-tuning and DPO post-training stages. We conjecture that this is because the role-play task in MT-Bench focuses more on functional roles rather than human-like characters, making it less sensitive to our boundary-aware learning approach.
@@
Table 5 exhibits the evaluation results on CharacterEval. It can be observed that Erabal significantly improves the baseline models’ performances on most metrics, especially on persona behavior (PB), role consistency (Cons.), and human likeness (HL). These metrics have high correlation to RPLAs’ general role-playing capabilities. DPO further improves the aforementioned three metrics, demonstrating its effectiveness.
@@
4.5 Human Assessment
To verify whether the model’s performance aligns with human evaluations, we conduct experiments using WikiRoleEval by replacing the LLM evaluators with human evaluators. Experimental results confirm that human assessments align with the automatic evaluations. For more details, see Appendix B.5.
5 Further Analysis
5.1 Effect of Data Scale
We conduct experiments based on LLaMA2-7B with 1/8, 1/4, 1/2, and full-sized datasets to study the impact of data scale. As depicted in Figure 4, as the data volume increased, almost all metrics showed rapid improvement. We believe that constructing more boundary-aware dialogues through Erabal will yield even better results. However, due to resource limitations, we limit our dataset to a maximum of approximately 16k dialogue data in this experiment. The performance on the Know. metric shows more significant improvement compared to the other three metrics, which is consistent with our previous analysis.
5.2 Effect of Model Scale
We further study the performance variation with Erabal on models of different scales. Experiments are conducted on LLaMA-7B, 13B, and 33B. Figure 5 shows the experimental results, from which we observe that the performance on the four metrics improves as the model size increases. However, for Cons. and Rej., the improvements tend to plateau when the model size exceeds 13B. Regarding role consistency and knowledge tasks, since these two tasks are relatively difficult, the performance improvements are more significant.